Suppose that the demand curve for wheat is Q 120-10p and the supply curve is Q-10p The government imposes a price ceiling of p $4 per unit. a. How do the equilibrium price and quantity change? (round quantities to the nearest integer and round prices to the nearest penny) The equilibrium quantity without the price ceiling is and the price without the price ceiling is S The equilibrium quantity with the price ceiling is b. What effect does this ceiling have on consumer surplus, producer surplus, and deadweight loss? The change in consumer surplus (CS) is S The change in producer surplus (PS) is S The deadweight loss (DWL) is S round your answer to the nearest penny). (round your answer to the nearest penny). (round your answer to the nearest penny).
Answer #1

Answer 1 : a : Equilibrium price and Quantity without price ceiling are :

Q= 120-10P ( Demand curve)

Q= 10P ( Supply curve)

120-10P= 10P

120= 20P

120/20= P

P= $6

Q= 10P= 10(6)= 60 units

The Equilibrium price without ceiling is $6 and Equilibrium Quantity without price ceiling is 60 units.

Equilibrium quantity with price ceiling

P= $4

Q= 10(4)= 40 units

b) The change in consumer surplus :

Consumer surplus without price ceiling = 0.5* ( Maximum price - Equilibrium price) * Equilibrium quantity= 0.5* (12-6)*60= $180

Consumer surplus with price ceiling= 0.5*( 12-4)*40= $160

Change in consumer surplus = $20

The change in Producer surplus :

Producer surplus without price ceiling= 0.5*( 6-0)*60= $180

Producer surplus with price ceiling =0.5* ( 4-0)*40= $80

Change in Producer surplus= $100

Dead weight loss=-0.5(4-6)*(40-60)=

-$20